### v1

# On the CP Nature of the ‘95 GeV’ Anomalies

Abstract
Under the assumption that the various evidences of a ‘95 GeV’ excess, seen in data at the Large Electron Positron (LEP) collider as well as the Large Hadron Collider (LHC), correspond to actual signals of new physics Beyond the Standard Model (BSM), we characterise the underlying particle explaining these in terms of its Charge/Parity (CP) quantum numbers in a model independent way. In doing so, we assume the new object having spin-0 and test the CP-even (scalar) and CP-odd (pseudoscalar) hypotheses as well superpositions of these in its $\tau^{+}\tau^{-}$ decays. We prove that the High Luminosity LHC (HL-LHC) will be in a position to disentangle the CP nature of such a new particle.
1 Introduction
Since the discovery of the 125 GeV Higgs boson at the Large Hadron Collider (LHC) in 2012 ATLAS:2012yve ; CMS:2012qbp , significant efforts have been made by the ATLAS and CMS collaborations in ascertaining the nature of this particle, that is, measuring its width, couplings, spin and CP quantum numbers. All corresponding measurements are pointing towards the discovered Higgs boson being the Standard Model (SM) one. Despite all this, though, the search for physics Beyond the SM (BSM) continues, spurred, on the theoretical side by its many flaws and, on the experimental side, by the ever increasing precision of the analyses as well as additional data sets being collected during Run 3 of the LHC.
The 2-Higgs Doublet Model (2HDM) Gunion:1992hs ; Branco:2011iw , as embedded in various BSM scenarios, including fundamental ones like Supersymmetry Moretti:2019ulc and Compositeness DeCurtis:2018zvh ; DeCurtis:2018iqd , is particularly interesting, as it offers in its particle spectrum both a CP-even and a CP-odd (neutral) Higgs boson, alongside the discovered SM-like Higgs state, both of which can in fact be lighter or heavier than the latter [^1] There are also charged Higgs bosons.. Hence, this BSM scenario has been studied extensively since its inception. However, its general version allows for non-diagonal Yukawa couplings, potentially leading to large Flavor Changing Neutral Currents (FCNCs) at tree level, contrary to experimental evidence. To constrain the latter, from the theoretical side, a $Z_{2}$ symmetry (which can be softly broken, in fact) is typically imposed to define the coupling structure of the two Higgs doublets to SM fermions. This classification includes the so-called Type-I, Type-II, lepton-specific and flipped scenarios Branco:2011iw , alongside the 2HDM Type-III, which allows direct couplings of both doublets to all SM fermions. The 2HDM Yukawa structure is then refined by both theoretical consistency requirements and experimental measurements of the $h$ mass and couplings or limits on the the same parameters for the $H,A$ and $H^{\pm}$ states.
Given that a 2HDM, as mentioned, allows (in some Types) for neutral Higgs bosons states ligther than 125 GeV, both ATLAS and CMS have pursued this possibility through dedicated searches in presence of a mass configuration wherein $m_{h}$ and/or $m_{A}<m_{H}=125$ GeV (with $m_{H^{\pm}}$ also rather light, to comply with various high-precision data). (Hereafter, in our notation, the CP-even $H$ state is the SM-like one already discovered while the $h$ and $A$ states are the (light) CP-even and CP-odd Higgs bosons, respectively, plus $H^{\pm}$ are the charged ones.)
[…]
Since all the excesses appear in very similar mass regions, several studies Cao:2016uwt ; Heinemeyer:2021msz ; Biekotter:2021qbc ; Biekotter:2019kde ; Cao:2019ofo ; Biekotter:2022abc ; Iguro:2022dok ; Li:2022etb ; Cline:2019okt ; Biekotter:2021ovi ; Crivellin:2017upt ; Cacciapaglia:2016tlr ; Abdelalim:2020xfk ; Biekotter:2022jyr ; Biekotter:2023jld ; Azevedo:2023zkg ; Biekotter:2023oen ; Cao:2024axg ; Wang:2024bkg ; Li:2023kbf ; Dev:2023kzu ; Borah:2023hqw ; Cao:2023gkc ; Ellwanger:2023zjc ; Aguilar-Saavedra:2023tql ; Ashanujjaman:2023etj ; Dutta:2023cig ; Ellwanger:2024txc ; Diaz:2024yfu ; Ellwanger:2024vvs ; Ayazi:2024fmn have explored the possibility of simultaneously explaining these anomalies within BSM frameworks featuring a non-standard Higgs state lighter than 125 GeV, while being in agreement with current measurements of the properties of the SM-like Higgs state observed at the LHC. In the attempt to explain the excesses in the $\gamma\gamma$ and $b\overline{b}$ channels, it was found in Refs. Benbrik:2022azi ; Benbrik:2022tlg ; Belyaev:2023xnv ; Belyaev:2024lah ; Benbrik:2024ptw that the 2HDM Type-III with a particular Yukawa texture can successfully accommodate both measurements simultaneously with the lightest CP-even Higgs boson of the model, while being consistent with all relevant theoretical and experimental constraints. In fact, also a 2HDM Type-I can afford similar explanations to the excesses Khanna:2024bah . Further recent studies have shown that actually all three aforementioned signatures can be simultaneously explained in the 2HDM plus a real (N2HDM) Biekotter:2022jyr or a complex (S2HDM) Biekotter:2023jld ; Biekotter:2023oen singlet.
The purpose of this paper is to assume that, indeed, a spin-0 particle of yet undefined CP is behind all such anomalies. However, amongst the latter, we focus here on the $\tau^{+}\tau^{-}$ one, as this has some diagnostic power in terms of pinning down its CP properties, as shown in Refs. Bower:2002zx ; Desch:2003rw ; Berge:2013jra ; Berge:2008wi ; Berge:2008dr ; Berge:2011ij ; Berge:2014sra ; Berge:2014wta ; Berge:2015nua ; Han:2016bvf ; Antusch:2020ngh ; Dutta:2021del ; Esmail:2024jdg . So, let us spend some time on this particular anomaly. The CMS collaboration has reported such an excess in the $\tau^{+}\tau^{-}$ final state across all $\tau$ decay modes (leptonic and hadronic) for both leptons. The local and global significances of this particular excess are $2.6\sigma$ and $2.3\sigma$ for $m_{\tau^{+}\tau^{-}}=95$ GeV, respectively. At $m_{\tau^{+}\tau^{-}}=100$ GeV, these values are $3.1\sigma$ and $2.7\sigma$. By introducing an additional single neutral narrow resonance $h$, the best fit values to the signal rate for this excess are CMS-PAS-HIG-21-001 :
$\sigma(gg\to h)\;{\rm BR}(h\to\tau^{+}\tau^{-})=7.7^{+3.9}_{-3.1}~{}\text{pb}\quad\text{for}~{}m_{h}=95~{}\text{GeV},$ (1)
$\sigma(gg\to h)\;{\rm BR}(h\to\tau^{+}\tau^{-})=5.8^{+2.4}_{-2.0}~{}\text{pb}\quad\text{for}~{}m_{h}=100~{}\text{GeV}.$ (2)
In our work, we assume a 95 GeV resonance. Interpetting Eq. (1) in terms of a signal strength, we get
$\mu^{\text{exp}}_{\tau^{+}\tau^{-}}=\frac{\sigma^{\text{exp}}(gg\to h\to\tau^{+}\tau^{-})}{\sigma^{\text{SM}}(gg\to H\to\tau^{+}\tau^{-})}=1.2\pm 0.5,$ (3)
where we use a symmetric uncertainty interval for such a signal strength, which is obtained from the lower uncertainty interval of the quoted result for $\sigma(gg\to h)\;{\rm BR}(h\to\tau^{+}\tau^{-})$. Also, in the signal strength, $H$ represents a fictitious SM Higgs boson with mass 95 GeV.
For completeness, we also report the signal strengths for the two other aforementioned anomalies, as follows:
$\mu^{\text{exp}}_{\gamma\gamma}=\frac{\sigma^{\text{exp}}(gg\to h\to\gamma\gamma)}{\sigma^{\text{SM}}(gg\to H\to\gamma\gamma)}=0.6\pm 0.2$ (4)
for $\gamma\gamma$ (see CMS:2018cyk ) and
$\mu^{\text{exp}}_{bb}=\frac{\sigma^{\text{exp}}(e^{+}e^{-}\to ZS\to Zb\bar{b})}{\sigma^{\text{SM}}(e^{+}e^{-}\to Zh\to Zb\bar{b})}=0.117\pm 0.057$ (5)
for $b\bar{b}$ (see LEPWorkingGroupforHiggsbosonsearches:2003ing ; ALEPH:2006tnd ).
In this work, in order to ascertain the CP state of the underlying spin-0 resonance (which, for illustrative purposes, we also assume to be a Higgs state), we have adopted a model agnostic explanation to all the excesses. As we particularly focus on the CP nature of this 95 GeV object, we have setup a field with Yukawa couplings to all fermions, i.e., proportional to their mass, scaled by the corresponding SM Yukawa, while allowing in such an interaction for the simultaneous presence of both a $\gamma_{5}$ (i.e., a pseudoscalar interaction) and a unit matrix (i.e., a scalar interaction) structure, in both the $ht\bar{t}$ and $h\tau^{+}\tau^{-}$ vertices, entering in production and decay, respectively. After exploring the parameter space of this simplified model by testing it against theoretical and experimental constraints, including fitting the described excesses, we look at unravelling the CP properties of the $h$ state at the LHC in the $\tau^{+}\tau^{-}$ channel using as diagnostic observable the signed acoplanarity angle $\phi_{\rm CP}$ between the two $\tau$ lepton decay planes. In order to reconstruct the latter, we have used one-prong hadronic $\tau$ decays and Impact Parameter (IP) information Berge:2014sra ; Berge:2015nua , wherein for the former we limited ourselves to the case of $\pi$ and $\rho$ mesons. By then performing a robust detector level simulation, we show that the it is indeed possible to unambiguously differentiate a CP-even from a CP-odd as well as a maximally CP-mixed signal at the LHC with 3000 fb-1 of data per experiment, which is the expected final dataset size of the High-Luminosity LHC (HL-LHC) Gianotti:2002xx ; Apollinari:2015wtw . However, we deem this result a conservative one since, as we gather more data at the current LHC (Run 3), the significance of the aforementioned excesses may well grow further, thereby stimulating dedicated analyses aimed at accessing the CP properties of this new resonance which may well prove to offer a better sensitivity than the one worked out in this article.
Our paper is organised as follows. In the next section, we describe our simplified model approach, including delineating its parameter space compatible with current constraints and explaining the $\gamma\gamma,\tau^{+}\tau^{-}$ and/or $b\bar{b}$ excesses. Then we move on to describe the construction of the mentioned CP-sensitive observable. Numerical results will follow, in turn preceding our conclusions.
2 A Simplified Model to Address the Excesses
We are interested in the study of CP properties of the observed spin-0 resonance, which might come from various UV complete scenarios like the Two Higgs Doublet Model, Next-to Minimal Supersymmetric scenarios (NMSSM) and other scalar extensions of the Standard Model. To remain model agnostic, we assumed a simplified model where, in addition to the SM spectrum, we consider only an additional scalar ($S$) of mass 95 GeV. The scalar particle $S$ couples with the fermions as follows:
$\mathcal{L}_{Sf\bar{f}}=-\rho^{S}_{f}\frac{m_{f}}{v}\Big{(}\cos\alpha\bar{f}f+i\sin\alpha\bar{f}\gamma_{5}f\Big{)}S.$ (6)
We have introduced CP-Violating (CPV) Yukawa coupling with $\alpha$ being the CPV parameter. $\alpha$ will be zero ($\pi/2$) when the scalar $S$ is a CP-even (CP-odd) particle. For simplicity, we have assumed a uniform $CP$ mixing among generations. The dimensionless real parameters $\rho^{S}_{f}$ ($f=t,b,\tau$) are the Yukawa coupling modifiers. The observed SM-like $CP$-even Higgs boson ($h$) of mass 125 GeV can couple with $S$ depending on the UV completion. Our analysis will not depend on the coupling of $S$ with the SM Higgs boson.
The coupling of the scalar $S$ with the weak neutral gauge boson, denoted as $g_{SVV}$, does not affect the LHC excess observed in $\tau^{+}\tau^{-}$ and $\gamma\gamma$ channel. However, a non-zero $g_{SVV}$ is necessary to explain the LEP $b\bar{b}$ excess. Such coupling can appear if $S$ is part of a weak multiplet which mixes with the SM Higgs doublet. Since the observed 125 GeV Higgs boson at the LHC closely resembles the SM, the mixing of the Higgs doublet can not be too large, and we have considered $g_{SVV}<0.3$.
2.1 Production and Decay Modes
Considering the large top mass and presuming $\rho^{h}_{b}$ not to be significantly large compared to $\rho^{h}_{t}$, we can obtain the gluon effective coupling described by a dimension five operator
$\mathcal{L}_{hgg}=\rho_{t}^{S}\cos\alpha~{}A_{S}~{}SG^{a}_{\mu\nu}G^{a,\mu\nu}+\rho_{t}^{S}\sin\alpha~{}A^{\prime}_{S}~{}SG^{a}_{\mu\nu}\widetilde{G}^{a,\mu\nu}$ (7)
where $G_{\mu\nu}$ is the gluon field strength tensor and $\widetilde{G}_{\mu\nu}=\frac{1}{4}\epsilon_{\mu\nu\rho\sigma}G^{\rho\sigma}$ is the corresponding dual field strength tensor and $A_{S}$ and $A^{\prime}_{S}$ are effective couplings. We incorporated Eq. (6) in FeynRules Christensen:2008py ; Alloul:2013bka and used SusHi Harlander:2012pb ; Harlander:2016hcx in order to estimate the NNLO production cross-section via gluon fusion. The gluon fusion cross section at the large top mass limit at NNLO is given by
$\sigma(pp\to gg\to S)={\rho_{t}^{S}}^{2}\Big{(}76.35\cos^{2}\alpha+176.32\sin^{2}\alpha\Big{)}~{}\text{pb},$ (8)
where the first term is the contribution from the $CP$ even part, and the second is the $CP$ odd contribution. The numerical coefficients are obtained from SusHi Harlander:2012pb ; Harlander:2016hcx .
For our analysis we have considered that the scalar $S$ can decay via the following channels Choi:2021nql ,
$S$$\to$ $f\bar{f},~{}VV^{*},~{}gg,~{}\gamma\gamma\textrm{ and }~{}Z\gamma.$ (9)
2.2 Allowed Parameter Space
Figure 1: The allowed parameter space for the $CP$ mixing angles $\alpha=0,~{}\pi/4$ and $\pi/2$. For the case of $\pi/2$, we cannot explain the LEP $b\bar{b}$ excess, and the allowed region is for the CMS $\tau^{+}\tau^{-}$ and $\gamma\gamma$ excesses. We can explain all the excesses for $\alpha=0$ and $\pi/4$.
In order to explain all the anomalies, the scalar $S$ must be coupled to third-generation fermions as well as massive gauge bosons. We fix the CPV parameter $\alpha$ and scan the free parameters $\rho_{f}^{S}$ to satisfy the excesses defined in Eq. (1), Eq. (4) and Eq. (5). We have shown the parameter space in Fig. 1, which can simultaneously explain all the excesses. For a CP-even scalar, the cross-section is relatively smaller, as shown in Eq. (8), and consequently, a large value of $\rho_{t}$ is required, as depicted in the top left panel of Fig. 1. For a mixed CP state, the required values of $\rho_{t}$ decreases as the cross-section increases due to a pseudoscalar component, the parameter is shown in the top-right panel of Fig. 1. In the bottom panel, we have considered a pure pseudoscalar model. Note that such a pure pseudoscalar scenario can not explain the LEP $b\bar{b}$ excess. Thus, the $g_{SVV}$ is redundant, and the parameter space does not depend on $g_{SVV}$. In all cases, the colour represents the $\rho_{b}$ value. From the figure, it is evident that it is possible to satisfy all three excesses simultaneously.
If we exclude the LEP excess, we can take $g_{SVV}=0$. Such a choice alters the decay spectrum of $S$ as $S\to Zff$ is not possible. In addition, it is possible to assume $\rho_{b}^{S}=0$ to explain the LHC excesses. Hence, to explain the LHC anomalies, as given in Eq. (1)-(4), a lower value of $\rho_{t/\tau}^{S}$ compared to the parameter space shown in Fig 1 will be enough, which corroborates Iguro:2022dok . Note that in Fig. 1, the parameter space does not depend on the sign of the angle $\alpha$, and thus, the results are identical for both ±$\pm\alpha$. As discussed in the next section, the CP observables will distinguish $+\alpha$ from $-\alpha$.
3 Uncovering the CP Nature of the Resonance
The CP transformation property of a scalar particle can be inferred from its decay to tau pairs since the correlation of the tau lepton polarisation planes is governed by the CP properties of the parent particle. From the reconstruction of the decay planes of the tau leptons, it is possible to infer the CP properties of the scalar particle.
The differential decay width of the scalar boson can be written as Kramer:1993jn ; Berge:2014sra ; Berge:2014wta
$d\Gamma_{S\to\tau^{+}\tau^{-}}\propto 1-\dfrac{\pi^{2}}{16}b(E_{+})b(E_{-})\cos(\phi_{CP}-2\alpha),$ (10)
where $b(E_{\pm})$ are the spectral function describing the spin analyzing power of a given decay mode Tsai:1971vv ; Berge:2014wta . For $\tau\to\pi\nu$ decay, $b(E_{\pm})=1$, and for other decay channels, it depends on the energy of the charged hadron at their respective tau-lepton rest frame. There are several methods  Bower:2002zx ; Desch:2003rw ; Berge:2013jra ; Berge:2008wi ; Berge:2008dr ; Berge:2011ij ; Berge:2014sra ; Berge:2014wta ; Berge:2015nua to estimate $\phi_{CP}$ and in this work we have used Impact Parameter (IP) method Berge:2014sra ; Berge:2015nua and the $\rho$ decay plane method  Bower:2002zx ; Berge:2015nua ; Han:2016bvf ; Dutta:2021del for hadronic $\tau$ decay. For completeness, we have described the methods briefly.
3.1 Impact Parameter Method
In the Impact Parameter (IP) method Berge:2014sra ; Berge:2015nua , the $\tau$-lepton decay plane is constructed using the 3-momentum vector ($\vec{q}^{\pm}$) and the three-dimensional (3D) impact parameter vector ($\vec{n}_{\pm}$) of the charged particles. The method can be used for any one-prong decay of the tau lepton, and in this work we have implemented the method for $S\to\tau^{+}\tau^{-}\to(\pi^{+}\nu_{\tau})(\pi^{-}\bar{\nu}_{\tau})$ decay. The IP vector in the laboratory frame is defined as the distance of the closest approach of the charged particle’s ($\pi^{\pm}$) track to the reconstructed primary vertex (PV) of an event. It is possible to reconstruct the 3D IP vector using the distance between the point of closest approach and the primary vertex in the $x-y$-plane ($d_{0}$) and the $z$-coordinate of the point ($z_{0}$ansen:2020ecn . We have used $d_{0}$ and $z_{0}$ and charged track momentum (highest $p_{T}$ charged track inside a $\tau$-tagged jet) provided by Delphes to find out the IP vector using the procedure describe in Hansen:2020ecn . The IP vector then converted to a four vector $n_{\pm}^{\mu}=(0,\vec{n}_{\pm})^{T}$ and boosted to zero-momentum frame (ZMF), the $\pi^{+}\pi^{-}$ rest frame. The boosted IP four-vectors are denoted as $n_{\pm}^{*\mu}$ and we define the angle,
$\phi^{*}=\arccos(\vec{n}^{*}_{+_{\perp}}.~{}\vec{n}^{*}_{-_{\perp}}),$ (11)
where $n^{*\mu}_{\pm_{\perp}}$ is the perpendicular component of $n^{*\mu}_{\pm}$ along the charged track $h^{\pm}$. To obtain a signed angle, we define the triple product, $\mathcal{O}(=q^{*-}\cdot(\vec{n}^{*}_{+_{\perp}}.~{}\vec{n}^{*}_{-_{\perp}}))$ where $q^{*-}$ is the 3-momentum vector of $\pi^{-}$ in ZMF. The CP observable is given by,
$\phi_{CP}=\left\{\begin{aligned} &\tilde{\phi}^{*}+\pi\quad&\mathcal{O}\geq 0\\ &\tilde{\phi}^{*}\quad&\mathcal{O}<0.\end{aligned}\right.$ (12)
3.2 Rho Decay Plane Method
When the $\tau$ lepton decays via the rho meson, $\tau\to\rho^{\pm}\nu_{\tau}\to\pi^{\pm}\pi^{0}\nu_{\tau}$, we use the information from the charged and neutral pion to define a CP-odd observable. In the $\pi^{+}\pi^{-}$ rest frame, the acoplanarity angle $\phi^{*}$ is defined between the plane spanned by the $(\pi^{+},\pi^{0})$ and $(\pi^{-},\pi^{0})$ vectors. The acoplanarity angle is given by Bower:2002zx ; Berge:2015nua ; Hansen:2020ecn ,
$\phi^{*}=\arccos(\vec{q}^{~{}*0}_{+_{\perp}}.~{}\vec{q}^{~{}*0}_{-_{\perp}}),$ (13)
where $\vec{q}^{~{}*0}_{+(-)_{\perp}}$ is the 3-momentum of the $\pi^{0}$ associated with $\pi^{+(-)}$ in the $\pi^{+}\pi^{-}$ rest-frame transverse to the direction of the associated charged pion. The signed angle $\tilde{\phi}^{*}$ is obtained using the CP-odd triple correlation $\mathcal{O_{CP}}(=q^{*-}\cdot(q^{*0}_{+_{\perp}}\times~{}q^{*0}_{-_{\perp}})$) and we define
$\tilde{\phi}^{*}=\left\{\begin{aligned} &\phi^{*}\quad&\mathcal{O_{CP}}\geq 0\\ &2\pi-\phi^{*}\quad&\mathcal{O_{CP}}<0.\end{aligned}\right.$ (14)
In the $S\to\tau^{+}\tau^{-}\to(\pi^{+}\pi^{0}\nu_{\tau})(\pi^{-}\pi^{0}\bar{\nu}_{\tau})$ decay, the spin analyzing power is proportional to $y_{\rho}=\dfrac{(E_{\pi^{+}}-E_{\pi^{0}})}{(E_{\pi^{+}}+E_{\pi^{0}})}\times\dfrac{(E_{\pi^{-}}-E_{\pi^{0}})}{(E_{\pi^{-}}+E_{\pi^{0}})}$ at the respective $\tau^{\pm}$ rest frame Bower:2002zx . Since $y_{\rho}$ can be both positive and negative and if we integrate over the charged and neutral pion moments, $\phi_{CP}$ does not display $CP$ sensitivity. We categorize the events depending on $y_{\rho}$ and define the final $CP$ sensitive variable,
$\phi_{CP}=\left\{\begin{aligned} &\tilde{\phi}^{*}+\pi\quad&y_{\rho}<0\\ &\tilde{\phi}^{*}\quad&\textrm{ otherwise}.\end{aligned}\right.$ (15)
To estimate $y_{\rho}$, we have calculated the energies of the pions in the lab frame, as it is not possible to reconstruct the $\tau$ rest frames. The choice of lab frame energy does not significantly reduce the asymmetry in $\phi_{CP}$ as long as the $\rho$ mesons are energetic ($p_{T}(\rho^{\pm})>20$ GeV) Berge:2015nua which is in general true for LHC.
3.3 Combined IP and Rho Decay Plane Methods
In the case of mixed decay, $\tau^{+}\tau^{-}\to(\pi\nu)(\rho\nu)$, we can combine the above-mentioned methods to define a CP-sensitive observable. This is possible if we consider the $\pi^{+}\pi^{-}$ rest frame for the $\rho$ Decay plane method.
In the case of $\tau^{+}\tau^{-}\to(\rho^{+}\nu_{\tau})(\pi^{-}\bar{\nu}_{\tau})$ decay, the angle is defined by Berge:2015nua ,
$\phi^{*}=\arccos(\vec{q}^{~{}*0}_{+_{\perp}}.~{}\vec{n}^{~{}*}_{-_{\perp}}),$ (16)
where, $\vec{q}^{~{}*}_{\pm}$ is the 3-momentum of $\pi^{\pm}$ and $\vec{n}^{~{}*}_{-}$ is the IP vector of $\pi^{-}$ and tall the vectors are in the $\pi^{+}\pi^{-}$ rest frame. As before, we can define a signed angle depending on the value of the triple product $\mathcal{O}_{3}(=\vec{q}^{~{}*}_{-}\cdot(\vec{q}^{~{}*0}_{+_{\perp}}\times~{}\vec{n}^{~{}*}_{-_{\perp}})$).
On the other hand, if $\tau^{+}$ decay to $\pi^{+}\nu_{\tau}$, we can define $\phi^{*}$ and the triple product appropriately. Note that, in both cases of mixed decay, we have considered the event classification based on $y_{\pm}(=(E_{\pi^{\pm}}-E_{\pi^{0}}))$.
3.4 Event Simulation and Selection
We implemented our simplified model for event simulation in FeynRules-2.3Alloul:2013bka . The signal and the background events are generated using MadGraph5-aMC@NLO-2.6.6Alwall:2011uj ; Alwall:2014hca . To handle the $\tau$ polarization, we have used the TauDecay Hagiwara:2012vz package in MadGraph5-aMC@NLO-2.6.6. For showering and hadronization we used PYTHIA-8.2 Sjostrand:2006za ; Sjostrand:2014zea , and detector effects were simulated using Delphes-3.4.2deFavereau:2013fsa . We use anti-kt algorithm Cacciari:2008gp for jet clustering with radius parameter $R=0.5$ and $p_{T}(j)>20$ GeV. For $\tau$-tagging we have used medium tag point with $70\%$ tagging and $5\times 10^{-3}$ mis-tagging efficiency. The signal events are classified as two $\tau$-tagged jets with $p_{T}>40$ GeV and $|\eta|<2.1$. We veto events with any $b$-tagged jets.
The background for such signal at the LHC is dominated by $Z\to\tau^{+}\tau^{-}$, $W^{\pm}+jets$ and QCD multi-jet events. We have generated sample background data to understand the nature of BG distribution for $CP$ sensitive observable. As we will discuss later, we have used actual data provided by the CMS collaboration through HEPDATA Maguire:2017ypu and rescaled it accordingly for different luminosity.
We obtain the charged and neutral pions from the $\tau$-jet constituents. From all the tracks within each $\tau$-jet, the track with the highest $p_{T}$ is considered to be the charged pion. Neutral pions deposit their energies in the hadron calorimeter, HCAL. From all the energy deposits, called towers, we estimate the one with the highest $E_{T}$ and assign the tower to the neutral pion. For the combination method, the charged pion track with higher $p_{T}$ is associated with the two-body decay of the $\tau$ ($\tau\to\pi\nu_{\tau}$).
4 Results
[…]
Now, we show how to uncover the CP properties of the scalar particle at the LHC. For a realistic representation of signal and background events, we have used the present CMS data of the $\tau^{+}\tau^{-}$ excess CMS:2018cyk ; CMS:2023yay . The data is taken from HEPDATAMaguire:2017ypu ; hepdata.128147 . For our analysis, we have selected the events for which the invariant mass $m_{\tau^{+}\tau^{-}}$ lies within $60-120$ GeV and we have selected the events with $p_{T}^{\tau^{+}\tau^{-}}>200$ GeV obtained by the vectorial $p_{T}$ sum of the visible decay products of tau lepton. This helps to minimise the background events significantly, as evident from Fig. 9 in ref CMS:2022goy .
To demonstrate the CP nature of the scalar, we plot the $\phi_{CP}$ for both signal and background events. The distribution peaks at a particular value of $\phi_{CP}$ for a given CP-admixture, whereas the background remains primarily flat. This enables estimating the CP nature from the $\phi_{CP}$ distribution alone. We found that with an integrated luminosity of 138 $fb^{-1}$, the total event distribution remains within statistical uncertainty of the background ($\sqrt{N_{B}}$). We considered high luminosity LHC (HL-LHC) Gianotti:2002xx ; Apollinari:2015wtw , and appropriately rescaled the existing data to an integrated luminosity of 3000 $fb^{-1}$. The $\phi_{CP}$ distribution for signal and background events for HL-LHC is shown in Fig. 2, where background events are shown in yellow with statistical uncertainty bars, and the signal events are shown in green. As anticipated, the background events exhibit a near-uniform distribution in all the cases. The top panel depicts the distribution for CP-Even (top-left) and CP-odd (top-right) scalar, whereas the bottom panels depict maximally CP-Mixed states. From all the figures, it is evident that for a given CP nature of the resonance, the cumulative number of signal and background events rise beyond the background uncertainty, signalling the CP property of the scalar.
$\alpha$ $\phi_{CP}$ Window $\mathcal{S}_{0}$ $\mathcal{S}_{\sqrt{b}}$ $\mathcal{S}_{1\%}$ $\mathcal{S}_{10\%}$
00 $120^{\circ}-240^{\circ}$ $3.10$ $2.18$ $2.01$ $0.26$
$\pi/2$ $0^{\circ}-80^{\circ}~{}~{}\&~{}~{}280^{\circ}-360^{\circ}$ $3.51$ $2.47$ $2.09$ $0.26$
$\pi/4$ $200^{\circ}-320^{\circ}$ $3.13$ $2.2$ $2.03$ $0.27$
-$\pi/4$ $40^{\circ}-160^{\circ}$ $3.12$ $2.2$ $2.03$ $0.27$
Table 1: Significance of different CP eigenstates with various background uncertainties at HL-LHC.
In order to quantify the result, we have chosen a window in $\phi_{CP}$ distribution and estimated a signal significance using the following formula Cowan:2010js ,
$\mathcal{S}_{\sigma_{b}}=\sqrt{2}\Big{\{}(N_{S}+N_{B})\ln\Big{[}\frac{(N_{S}+N_{B})(N_{B}+\sigma_{b}^{2})}{N_{B}^{2}+(N_{S}+N_{B})\sigma_{b}^{2}}\Big{]}-\frac{N_{B}^{2}}{\sigma_{b}^{2}}\ln\Big{[}1+\frac{\sigma^{2}_{B}N_{S}}{N_{B}(N_{B}+\sigma_{b}^{2})}\Big{]}\Big{\}}^{1/2},$ (17)
where $N_{S(B)}$ denotes signal(background) events and $\sigma_{b}$ denotes the background uncertainty. In Tab. 1 we have tabulated the significance for an optimal $\phi_{CP}$ window with various background uncertainty to achieve a robust estimation of the Significance. As expected, vanishing BG uncertainty provides a significance of more than $3~{}\sigma$ at HL-LHC. As uncertainty increases, Significance reduces but remains large enough if we consider the background uncertainty solely statistical ($\sqrt{N_{B}}$) or $1\%$. If $\sigma_{b}$ becomes large ($\sim 10\%$), the number of signal events need to be higher to achieve any significance.
5 Conclusions
In summary, under the assumption that several anomalies seen in the invariant mass of $\gamma\gamma$, $\tau^{+}\tau^{-}$ (from LHC data) and $b\bar{b}$ (from LEP data) events are due to a single spin-0 object with mass 95 GeV having Yukawa couplings to third generation fermions, we have performed a detector level analysis aimed at extracting the CP properties of such an object. We have done so limitedly to the $\tau^{+}\tau^{-}$ final state, as it offers the possibility of accessing the corresponding CP quantum numbers whichever the $\tau$ decay, whether leptonic or hadronic, through the relative angle between the two decay planes defined through the visible particles emerging from the $\tau$ lepton decays. Our approach has been based on a simplified model, which is characterized by the relevant Yukawa interactions being defined in terms of correcting factors to the corresponding SM vertex strengths while allowing for both a scalar and pseudoscalar interaction as well as any superposition thereof.
We have found that, under the assumption that the $\gamma\gamma$ and $\tau^{+}\tau^{-}$ anomalies will persist throughout the LHC era, the HL-LHC, in standard energy and luminosity configurations, will offer the chance to extract the CP properties of such a new spin-0 resonance. Once the latter is combined with cross-section and BR information, it would then be possible to map the results obtained in our model-independent parameterization into well-defined models like, e.g., a 2HDM of whichever Type, so long that it can host a suitable 95 GeV state in its Higgs particle spectrum, as it has indeed been proven to be the case in previous literature.
We therefore advocate experimental analyses by ATLAS and CMS aimed at implementing our approach, or indeed improved versions of it, with upcoming datasets at the CERN machine.
Acknowledgements
T.M. and P.S. would like to thank Artur Gottmann and Ravindra Kumar Verma for some useful discussions. T.M. is supported by the BITS Pilani Grant NFSG/PIL/2023/P3801. S.M. is supported in part through the NExT Institute and STFC Consolidated Grant No. ST/L000296/1.

### v2

@@
 […]
-Since all the excesses appear in very similar mass regions, several studies Cao:2016uwt ; Heinemeyer:2021msz ; Biekotter:2021qbc ; Biekotter:2019kde ; Cao:2019ofo ; Biekotter:2022abc ; Iguro:2022dok ; Li:2022etb ; Cline:2019okt ; Biekotter:2021ovi ; Crivellin:2017upt ; Cacciapaglia:2016tlr ; Abdelalim:2020xfk ; Biekotter:2022jyr ; Biekotter:2023jld ; Azevedo:2023zkg ; Biekotter:2023oen ; Cao:2024axg ; Wang:2024bkg ; Li:2023kbf ; Dev:2023kzu ; Borah:2023hqw ; Cao:2023gkc ; Ellwanger:2023zjc ; Aguilar-Saavedra:2023tql ; Ashanujjaman:2023etj ; Dutta:2023cig ; Ellwanger:2024txc ; Diaz:2024yfu ; Ellwanger:2024vvs ; Ayazi:2024fmn have explored the possibility of simultaneously explaining these anomalies within BSM frameworks featuring a non-standard Higgs state lighter than 125 GeV, while being in agreement with current measurements of the properties of the SM-like Higgs state observed at the LHC.
+Since all the excesses appear in very similar mass regions, several studies Cao:2016uwt ; Heinemeyer:2021msz ; Biekotter:2021qbc ; Biekotter:2019kde ; Cao:2019ofo ; Biekotter:2022abc ; Iguro:2022dok ; Li:2022etb ; Cline:2019okt ; Biekotter:2021ovi ; Crivellin:2017upt ; Cacciapaglia:2016tlr ; Abdelalim:2020xfk ; Biekotter:2022jyr ; Biekotter:2023jld ; Azevedo:2023zkg ; Biekotter:2023oen ; Cao:2024axg ; Wang:2024bkg ; Li:2023kbf ; Dev:2023kzu ; Borah:2023hqw ; Cao:2023gkc ; Ellwanger:2023zjc ; Aguilar-Saavedra:2023tql ; Ashanujjaman:2023etj ; Bhattacharya:2023lmu ; Dutta:2023cig ; Ellwanger:2024txc ; Diaz:2024yfu ; Ellwanger:2024vvs ; Ayazi:2024fmn ; Arhrib:2024wjj ; Gao:2024qag have explored the possibility of simultaneously explaining these anomalies within BSM frameworks featuring a non-standard Higgs state lighter than 125 GeV, while being in agreement with current measurements of the properties of the SM-like Higgs state observed at the LHC.
 In the attempt to explain the excesses in the $\gamma\gamma$ and $b\overline{b}$ channels, it was found in Refs.
@@
 We implemented our simplified model for event simulation in FeynRules-2.3Alloul:2013bka .
-The signal and the background events are generated using MadGraph5-aMC@NLO-2.6.6Alwall:2011uj ; Alwall:2014hca .
-To handle the $\tau$ polarization, we have used the TauDecay Hagiwara:2012vz package in MadGraph5-aMC@NLO-2.6.6.
-For showering and hadronization we used PYTHIA-8.2 Sjostrand:2006za ; Sjostrand:2014zea , and detector effects were simulated using Delphes-3.4.2deFavereau:2013fsa .
+The signal and the background events are generated using MadGraph5-aMC@NLO-3.5.3Alwall:2011uj ; Alwall:2014hca .
+To handle the $\tau$ polarization, we have used the TauDecay Hagiwara:2012vz package in MadGraph5-aMC@NLO-3.5.3.
+For showering and hadronization we used PYTHIA-8.3 Sjostrand:2014zea ; Bierlich:2022pfr , and detector effects were simulated using Delphes-3.5.0deFavereau:2013fsa .
 We use anti-kt algorithm Cacciari:2008gp for jet clustering with radius parameter $R=0.5$ and $p_{T}(j)>20$ GeV.
@@
 From all the tracks within each $\tau$-jet, the track with the highest $p_{T}$ is considered to be the charged pion.
-Neutral pions deposit their energies in the hadron calorimeter, HCAL.
+Neutral pions decay promptly into a pair of photons and therefore deposit their energies in the electromagnetic calorimeter, ECAL.
 From all the energy deposits, called towers, we estimate the one with the highest $E_{T}$ and assign the tower to the neutral pion.

### v3

@@
 Abstract
-Under the assumption that the various evidences of a ‘95 GeV’ excess, seen in data at the Large Electron Positron (LEP) collider as well as the Large Hadron Collider (LHC), correspond to actual signals of new physics Beyond the Standard Model (BSM), we characterise the underlying particle explaining these in terms of its Charge/Parity (CP) quantum numbers in a model independent way.
-In doing so, we assume the new object having spin-0 and test the CP-even (scalar) and CP-odd (pseudoscalar) hypotheses as well superpositions of these in its $\tau^{+}\tau^{-}$ decays.
-We prove that the High Luminosity LHC (HL-LHC) will be in a position to disentangle the CP nature of such a new particle.
+Under the assumption that the various evidences of a ‘95 GeV’ excess, seen in data at the Large Electron Positron (LEP) collider as well as the Large Hadron Collider (LHC), correspond to actual signals of new physics Beyond the Standard Model (BSM), we characterise the underlying particle explaining these anomalies in terms of its Charge/Parity (CP) quantum numbers.
+In doing so, we use $\chi^{2}$ fits to test the CP-even (scalar) and CP-odd (pseudoscalar) hypotheses and superpositions of these, thus under the assumption of a spin-0 resonance.
+This is done through the exploitation of $\tau^{+}\tau^{-}$ decays, in both their fully hadronic and semi-leptonic modes, in a model-independent way, so that our approach enables one to test a variety of BSM hypotheses, having proven here that the High-Luminosity LHC (HL-LHC) will be in a position to disentangle the CP nature of such a new particle within $\pm(0.27-0.47)$ radians of the true hypothesis at $90\%$ Confidence Level (CL), depending on the assumed background systematics.
+preprint: CQUeST-2025-0763
 1 Introduction
@@
 All corresponding measurements are pointing towards the discovered Higgs boson being the Standard Model (SM) one.
-Despite all this, though, the search for physics Beyond the SM (BSM) continues, spurred, on the theoretical side by its many flaws and, on the experimental side, by the ever increasing precision of the analyses as well as additional data sets being collected during Run 3 of the LHC.
-The 2-Higgs Doublet Model (2HDM) Gunion:1992hs ; Branco:2011iw , as embedded in various BSM scenarios, including fundamental ones like Supersymmetry Moretti:2019ulc and Compositeness DeCurtis:2018zvh ; DeCurtis:2018iqd , is particularly interesting, as it offers in its particle spectrum both a CP-even and a CP-odd (neutral) Higgs boson, alongside the discovered SM-like Higgs state, both of which can in fact be lighter or heavier than the latter [^1] There are also charged Higgs bosons..
+Despite all this, though, the search for physics Beyond the SM (BSM) continues, spurred, on the theoretical side, by its many flaws and, on the experimental side, by the ever increasing precision of the analyses as well as additional data sets being collected during Run 3 of the LHC.
+The 2-Higgs Doublet Model (2HDM) Gunion:1992hs ; Branco:2011iw , as embedded in various BSM scenarios, including fundamental ones like Supersymmetry Moretti:2019ulc and Compositeness DeCurtis:2018zvh ; DeCurtis:2018iqd , is particularly interesting, as it offers in its particle spectrum both a CP-even and a CP-odd (neutral) Higgs boson, alongside the discovered SM-like Higgs state, both of which can in fact be lighter or heavier than the latter [^1] The CP-even (neutral) Higgs states of the 2HDM are labelled as $h$ and $H$ (with $m_{h}<m_{H}$) while the CP-odd (neutral) one as $A$, plus there are also two charged Higgs bosons, $H^{\pm}$..
 Hence, this BSM scenario has been studied extensively since its inception.
@@
 This classification includes the so-called Type-I, Type-II, lepton-specific and flipped scenarios Branco:2011iw , alongside the 2HDM Type-III, which allows direct couplings of both doublets to all SM fermions.
-The 2HDM Yukawa structure is then refined by both theoretical consistency requirements and experimental measurements of the $h$ mass and couplings or limits on the the same parameters for the $H,A$ and $H^{\pm}$ states.
-Given that a 2HDM, as mentioned, allows (in some Types) for neutral Higgs bosons states ligther than 125 GeV, both ATLAS and CMS have pursued this possibility through dedicated searches in presence of a mass configuration wherein $m_{h}$ and/or $m_{A}<m_{H}=125$ GeV (with $m_{H^{\pm}}$ also rather light, to comply with various high-precision data).
-(Hereafter, in our notation, the CP-even $H$ state is the SM-like one already discovered while the $h$ and $A$ states are the (light) CP-even and CP-odd Higgs bosons, respectively, plus $H^{\pm}$ are the charged ones.)
+The 2HDM Yukawa structure is then refined by both theoretical consistency requirements and experimental measurements of the SM-like Higgs mass and couplings or limits on the same parameters for the other Higgs states.
+Given that a 2HDM, as mentioned, allows (in some Types) for neutral Higgs bosons states lighter than 125 GeV, both ATLAS and CMS have pursued this possibility through dedicated searches in presence of a mass configuration wherein $m_{h}$ and/or $m_{A}<m_{H}=125$ GeV (with $m_{H^{\pm}}$ also rather small, to comply with various high-precision data).
+(Hereafter, in our notation, when referring to the 2HDM, the CP-even $H$ state is the SM-like one already discovered while the $h$ and $A$ states are the lightest CP-even and CP-odd Higgs bosons, respectively.) However, in performing such searches, neither ATLAS nor CMS have defined potential signal regions based on 2HDM assumptions, as therein the latter only entered at the interpretation level.
+Therefore, their results can be used in a model-independent way, which is what we will be doing here by introducing a simplified model approach wherein the underlying BSM scenario is parameterized solely in terms of masses and (gauge plus Yukawa) couplings that are fitted to data, yet, we assume the same particle content of the generic 2HDM when computing the decay width of a new Higgs state that we will introduce (hereafter, denoted by $S$).
 […]
-Since all the excesses appear in very similar mass regions, several studies Cao:2016uwt ; Heinemeyer:2021msz ; Biekotter:2021qbc ; Biekotter:2019kde ; Cao:2019ofo ; Biekotter:2022abc ; Iguro:2022dok ; Li:2022etb ; Cline:2019okt ; Biekotter:2021ovi ; Crivellin:2017upt ; Cacciapaglia:2016tlr ; Abdelalim:2020xfk ; Biekotter:2022jyr ; Biekotter:2023jld ; Azevedo:2023zkg ; Biekotter:2023oen ; Cao:2024axg ; Wang:2024bkg ; Li:2023kbf ; Dev:2023kzu ; Borah:2023hqw ; Cao:2023gkc ; Ellwanger:2023zjc ; Aguilar-Saavedra:2023tql ; Ashanujjaman:2023etj ; Bhattacharya:2023lmu ; Dutta:2023cig ; Ellwanger:2024txc ; Diaz:2024yfu ; Ellwanger:2024vvs ; Ayazi:2024fmn ; Arhrib:2024wjj ; Gao:2024qag have explored the possibility of simultaneously explaining these anomalies within BSM frameworks featuring a non-standard Higgs state lighter than 125 GeV, while being in agreement with current measurements of the properties of the SM-like Higgs state observed at the LHC.
+Since all the excesses appear in very similar mass regions, several studies Cao:2016uwt ; Heinemeyer:2021msz ; Biekotter:2021qbc ; Biekotter:2019kde ; Cao:2019ofo ; Biekotter:2022abc ; Iguro:2022dok ; Li:2022etb ; Cline:2019okt ; Biekotter:2021ovi ; Crivellin:2017upt ; Cacciapaglia:2016tlr ; Abdelalim:2020xfk ; Biekotter:2022jyr ; Biekotter:2023jld ; Azevedo:2023zkg ; Biekotter:2023oen ; Cao:2024axg ; Wang:2024bkg ; Li:2023kbf ; Dev:2023kzu ; Borah:2023hqw ; Cao:2023gkc ; Ellwanger:2023zjc ; Aguilar-Saavedra:2023tql ; Ashanujjaman:2023etj ; Bhattacharya:2023lmu ; Dutta:2023cig ; Ellwanger:2024txc ; Diaz:2024yfu ; Ellwanger:2024vvs ; Ayazi:2024fmn ; Arhrib:2024wjj ; Gao:2024qag have explored the possibility of simultaneously explaining these anomalies within BSM frameworks featuring a Higgs state lighter than 125 GeV, while being in agreement with current measurements of the properties of the SM-like Higgs state observed at the LHC.
 In the attempt to explain the excesses in the $\gamma\gamma$ and $b\overline{b}$ channels, it was found in Refs.
@@
 At $m_{\tau^{+}\tau^{-}}=100$ GeV, these values are $3.1\sigma$ and $2.7\sigma$.
-By introducing an additional single neutral narrow resonance $h$, the best fit values to the signal rate for this excess are CMS-PAS-HIG-21-001 :
-$\sigma(gg\to h)\;{\rm BR}(h\to\tau^{+}\tau^{-})=7.7^{+3.9}_{-3.1}~{}\text{pb}\quad\text{for}~{}m_{h}=95~{}\text{GeV},$ (1)
-$\sigma(gg\to h)\;{\rm BR}(h\to\tau^{+}\tau^{-})=5.8^{+2.4}_{-2.0}~{}\text{pb}\quad\text{for}~{}m_{h}=100~{}\text{GeV}.$ (2)
-In our work, we assume a 95 GeV resonance. Interpetting Eq. (1) in terms of a signal strength, we get
-$\mu^{\text{exp}}_{\tau^{+}\tau^{-}}=\frac{\sigma^{\text{exp}}(gg\to h\to\tau^{+}\tau^{-})}{\sigma^{\text{SM}}(gg\to H\to\tau^{+}\tau^{-})}=1.2\pm 0.5,$ (3)
-where we use a symmetric uncertainty interval for such a signal strength, which is obtained from the lower uncertainty interval of the quoted result for $\sigma(gg\to h)\;{\rm BR}(h\to\tau^{+}\tau^{-})$.
-Also, in the signal strength, $H$ represents a fictitious SM Higgs boson with mass 95 GeV.
+By introducing an additional single neutral narrow resonance $S$, the best fit values to the signal rate for this excess are CMS-PAS-HIG-21-001 :
+$\sigma(gg\to S)\;{\rm BR}(S\to\tau^{+}\tau^{-})=7.7^{+3.9}_{-3.1}~\text{pb}\quad\text{for}~m_{S}=95~\text{GeV},$ (1)
+$\sigma(gg\to S)\;{\rm BR}(S\to\tau^{+}\tau^{-})=5.8^{+2.4}_{-2.0}~\text{pb}\quad\text{for}~m_{S}=100~\text{GeV}.$ (2)
+In our work, we assume a 95 GeV resonance. Interpreting Eq. (1) in terms of a signal strength, we get
+$\mu^{\text{exp}}_{\tau^{+}\tau^{-}}=\frac{\sigma^{\text{exp}}(gg\to S\to\tau^{+}\tau^{-})}{\sigma^{\text{SM}}(gg\to\Phi\to\tau^{+}\tau^{-})}=1.2\pm 0.5,$ (3)
+where we use a symmetric uncertainty interval for such a signal strength, which is obtained from the lower uncertainty interval of the quoted result for $\sigma(gg\to S)\;\times\;{\rm BR}(S\to\tau^{+}\tau^{-})$.
+Also, in the signal strength, $\Phi$ represents a fictitious SM Higgs boson with mass 95 GeV.
 For completeness, we also report the signal strengths for the two other aforementioned anomalies, as follows:
-$\mu^{\text{exp}}_{\gamma\gamma}=\frac{\sigma^{\text{exp}}(gg\to h\to\gamma\gamma)}{\sigma^{\text{SM}}(gg\to H\to\gamma\gamma)}=0.6\pm 0.2$ (4)
+$\mu^{\text{exp}}_{\gamma\gamma}=\frac{\sigma^{\text{exp}}(gg\to S\to\gamma\gamma)}{\sigma^{\text{SM}}(gg\to\Phi\to\gamma\gamma)}=0.6\pm 0.2$ (4)
 for $\gamma\gamma$ (see CMS:2018cyk ) and
-$\mu^{\text{exp}}_{bb}=\frac{\sigma^{\text{exp}}(e^{+}e^{-}\to ZS\to Zb\bar{b})}{\sigma^{\text{SM}}(e^{+}e^{-}\to Zh\to Zb\bar{b})}=0.117\pm 0.057$ (5)
+$\mu^{\text{exp}}_{bb}=\frac{\sigma^{\text{exp}}(e^{+}e^{-}\to ZS\to Zb\bar{b})}{\sigma^{\text{SM}}(e^{+}e^{-}\to Z\Phi\to Zb\bar{b})}=0.117\pm 0.057$ (5)
 for $b\bar{b}$ (see LEPWorkingGroupforHiggsbosonsearches:2003ing ; ALEPH:2006tnd ).
-In this work, in order to ascertain the CP state of the underlying spin-0 resonance (which, for illustrative purposes, we also assume to be a Higgs state), we have adopted a model agnostic explanation to all the excesses.
-As we particularly focus on the CP nature of this 95 GeV object, we have setup a field with Yukawa couplings to all fermions, i.e., proportional to their mass, scaled by the corresponding SM Yukawa, while allowing in such an interaction for the simultaneous presence of both a $\gamma_{5}$ (i.e., a pseudoscalar interaction) and a unit matrix (i.e., a scalar interaction) structure, in both the $ht\bar{t}$ and $h\tau^{+}\tau^{-}$ vertices, entering in production and decay, respectively.
-After exploring the parameter space of this simplified model by testing it against theoretical and experimental constraints, including fitting the described excesses, we look at unravelling the CP properties of the $h$ state at the LHC in the $\tau^{+}\tau^{-}$ channel using as diagnostic observable the signed acoplanarity angle $\phi_{\rm CP}$ between the two $\tau$ lepton decay planes.
-In order to reconstruct the latter, we have used one-prong hadronic $\tau$ decays and Impact Parameter (IP) information Berge:2014sra ; Berge:2015nua , wherein for the former we limited ourselves to the case of $\pi$ and $\rho$ mesons.
-By then performing a robust detector level simulation, we show that the it is indeed possible to unambiguously differentiate a CP-even from a CP-odd as well as a maximally CP-mixed signal at the LHC with 3000 fb-1 of data per experiment, which is the expected final dataset size of the High-Luminosity LHC (HL-LHC) Gianotti:2002xx ; Apollinari:2015wtw .
-However, we deem this result a conservative one since, as we gather more data at the current LHC (Run 3), the significance of the aforementioned excesses may well grow further, thereby stimulating dedicated analyses aimed at accessing the CP properties of this new resonance which may well prove to offer a better sensitivity than the one worked out in this article.
+In this work, in order to ascertain the CP state of the underlying spin-0 resonance (which, for illustrative purposes, we also assume to be a Higgs state), we have adopted a model-agnostic explanation to all the excesses.
+As we particularly focus on the CP nature of this 95 GeV object, we have setup a field with Yukawa couplings to the third generation fermions, scaled by the corresponding SM Yukawa, while allowing in such an interaction for the simultaneous presence of both a $\gamma_{5}$ (i.e., a pseudoscalar interaction) and a unit matrix (i.e., a scalar interaction) structure, in both the $St\bar{t}$ and $S\tau^{+}\tau^{-}$ vertices, entering in production and decay, respectively.
+After exploring the parameter space of this simplified model by testing it against theoretical and experimental constraints, including fitting the described excesses, we look at unravelling the CP properties of the $S$ state at the LHC in the $\tau^{+}\tau^{-}$ channel using as diagnostic observable the signed acoplanarity angle $\phi_{\rm CP}$ between the two $\tau$ lepton decay planes.
+In order to reconstruct the decay planes, we have used both hadronic and semi-leptonic decays of the tau lepton, and used the Impact Parameter (IP) information Berge:2014sra ; Berge:2015nua of the charged prong as well as the $\rho$-decay method Bower:2002zx ; Berge:2015nua ; Hansen:2020ecn .
+By then performing a robust detector-level simulation, we show that it is indeed possible to unambiguously differentiate a CP-even from a CP-odd as well as a maximally CP-mixed signal at the LHC with 3000 fb-1 of data per experiment, which is the expected final dataset size of the HL-LHC Gianotti:2002xx ; Apollinari:2015wtw .
+However, we deem this result a conservative one since, as we gather more data at the current LHC (Run 3), the significance of the aforementioned excesses may well grow further, thereby stimulating especially dedicated analyses aimed at accessing the CP properties of this potential new resonance, which may well prove to offer an even better sensitivity than the one established in this article.
+Yet, our findings are altogether already very encouraging in this respect.
+For the observed 125 GeV scalar, a similar CP characterization has been done by ATLAS ATLAS:2022akr and CMS collaboration CMS:2021sdq .
+However, at the lower mass, the signal rates, backgrounds as well as kinematics are different, and no dedicated CP study exists in this mass region.
+We believe this is important in order to test various hypotheses made in the literature about the CP properties of possible interpretations.
 Our paper is organised as follows.
@@
 2 A Simplified Model to Address the Excesses
-We are interested in the study of CP properties of the observed spin-0 resonance, which might come from various UV complete scenarios like the Two Higgs Doublet Model, Next-to Minimal Supersymmetric scenarios (NMSSM) and other scalar extensions of the Standard Model.
+We are interested in the study of CP properties of the observed spin-0 resonance, which might come from various UV complete scenarios like the 2HDM, Next-to Minimal Supersymmetric Standard Model (NMSSM) and other scalar extensions of the SM.
 To remain model agnostic, we assumed a simplified model where, in addition to the SM spectrum, we consider only an additional scalar ($S$) of mass 95 GeV.
-The scalar particle $S$ couples with the fermions as follows:
+The scalar particle $S$ couples with the third generation fermions as follows:
 $\mathcal{L}_{Sf\bar{f}}=-\rho^{S}_{f}\frac{m_{f}}{v}\Big{(}\cos\alpha\bar{f}f+i\sin\alpha\bar{f}\gamma_{5}f\Big{)}S.$ (6)
 We have introduced CP-Violating (CPV) Yukawa coupling with $\alpha$ being the CPV parameter.
-$\alpha$ will be zero ($\pi/2$) when the scalar $S$ is a CP-even (CP-odd) particle.
+Here, $\alpha$ will be zero ($\pi/2$) when the scalar $S$ is a CP-even (CP-odd) Higgs particle (e.g., in the 2HDM, this will be $H$ (AA)).
 For simplicity, we have assumed a uniform CP mixing among generations.
 The dimensionless real parameters $\rho^{S}_{f}$ ($f=t,b,\tau$) are the Yukawa coupling modifiers.
-The observed SM-like $CP$-even Higgs boson ($h$) of mass 125 GeV can couple with $S$ depending on the UV completion.
+The observed SM-like CP-even Higgs boson of mass 125 GeV can couple with $S$ depending on the Ultra-Violet (UV) completion.
 Our analysis will not depend on the coupling of $S$ with the SM Higgs boson.
-The coupling of the scalar $S$ with the weak neutral gauge boson, denoted as $g_{SVV}$, does not affect the LHC excess observed in $\tau^{+}\tau^{-}$ and $\gamma\gamma$ channel.
+The coupling of the scalar $S$ with the weak neutral gauge boson, denoted as $g_{SVV}$, does not affect the LHC excess observed in the $\tau^{+}\tau^{-}$ and $\gamma\gamma$ channels.
 However, a non-zero $g_{SVV}$ is necessary to explain the LEP $b\bar{b}$ excess.
 Such coupling can appear if $S$ is part of a weak multiplet which mixes with the SM Higgs doublet.
-Since the observed 125 GeV Higgs boson at the LHC closely resembles the SM, the mixing of the Higgs doublet can not be too large, and we have considered $g_{SVV}<0.3$.
+Since the observed 125 GeV Higgs boson at the LHC closely resembles the SM one, the mixing of the Higgs doublet cannot be too large, and we have considered $g_{SVV}<0.3$.
 2.1 Production and Decay Modes
-Considering the large top mass and presuming $\rho^{h}_{b}$ not to be significantly large compared to $\rho^{h}_{t}$, we can obtain the gluon effective coupling described by a dimension five operator
+Considering the large top mass and presuming $\rho^{S}_{b}$ not to be significantly large compared to $\rho^{S}_{t}$, we can obtain the gluon effective coupling described by a dimension five operator
 $\mathcal{L}_{hgg}=\rho_{t}^{S}\cos\alpha~A_{S}~SG^{a}_{\mu\nu}G^{a,\mu\nu}+\rho_{t}^{S}\sin\alpha~A^{\prime}_{S}~SG^{a}_{\mu\nu}\widetilde{G}^{a,\mu\nu}$ (7)
@@
 We incorporated Eq.
-(6) in FeynRules Christensen:2008py ; Alloul:2013bka and used SusHi Harlander:2012pb ; Harlander:2016hcx in order to estimate the NNLO production cross-section via gluon fusion.
-The gluon fusion cross section at the large top mass limit at NNLO is given by
-$\sigma(pp\to gg\to S)={\rho_{t}^{S}}^{2}\Big{(}76.35\cos^{2}\alpha+176.32\sin^{2}\alpha\Big{)}~{}\text{pb},$ (8)
-where the first term is the contribution from the $CP$ even part, and the second is the $CP$ odd contribution. The numerical coefficients are obtained from SusHi Harlander:2012pb ; Harlander:2016hcx .
-For our analysis we have considered that the scalar $S$ can decay via the following channels Choi:2021nql ,
+(6) in FeynRules Christensen:2008py ; Alloul:2013bka and used SusHi Harlander:2012pb ; Harlander:2016hcx in order to estimate the Next-to-Next-to-Leading Order (NNLO) production cross section via gluon fusion.
+The gluon fusion cross section in the large top mass limit at NNLO is given by
+$\sigma(gg\to S)={\rho_{t}^{S}}^{2}\Big{(}76.35\cos^{2}\alpha+176.32\sin^{2}\alpha\Big{)}~\text{pb},$ (8)
+where the first term is the contribution from the CP-even part, and the second is the CP-odd contribution. The numerical coefficients are obtained from SusHi Harlander:2012pb ; Harlander:2016hcx .
+Finally, for our analysis, we have considered that the scalar $S$ can decay via the following channels Choi:2021nql :
 $S$ $\to$ $f\bar{f},~VV^{*},~gg,~\gamma\gamma\textrm{ and }~Z\gamma.$ (9)
@@
 (5).
-We have shown the parameter space in Fig.
-1, which can simultaneously explain all the excesses.
-For a CP-even scalar, the cross-section is relatively smaller, as shown in Eq.
-(8), and consequently, a large value of $\rho_{t}$ is required, as depicted in the top left panel of Fig.
+We have shown the parameter space that can simultaneously explain all the excesses in Fig.
 1.
-For a mixed CP state, the required values of $\rho_{t}$ decreases as the cross-section increases due to a pseudoscalar component, the parameter is shown in the top-right panel of Fig.
+For a CP-even scalar, the cross section is relatively small, as shown in Eq.
+(8), and consequently, a large value of $\rho_{t}$ is required, as depicted in the top-left panel of Fig.
+1.
+For a CP-mixed state, the required values of $\rho_{t}$ decrease as the cross section increases due to a pseudoscalar component, the parameter is shown in the top-right panel of Fig.
 1.
@@
 Note that such a pure pseudoscalar scenario can not explain the LEP $b\bar{b}$ excess.
-Thus, the $g_{SVV}$ is redundant, and the parameter space does not depend on $g_{SVV}$.
-In all cases, the colour represents the $\rho_{b}$ value.
+Thus, the $g_{SVV}$ is redundant, and the viable parameter space does not depend on it.
+In all cases, the color represents the $\rho_{b}$ value.
 From the figure, it is evident that it is possible to satisfy all three excesses simultaneously.
@@
 In addition, it is possible to assume $\rho_{b}^{S}=0$ to explain the LHC excesses.
-Hence, to explain the LHC anomalies, as given in Eq.
-(1)-(4), a lower value of $\rho_{t/\tau}^{S}$ compared to the parameter space shown in Fig 1 will be enough, which corroborates Iguro:2022dok .
-Note that in Fig.
-1, the parameter space does not depend on the sign of the angle $\alpha$, and thus, the results are identical for both ±$\pm\alpha$.
-As discussed in the next section, the CP observables will distinguish $+\alpha$ from $-\alpha$.
+Hence, to explain the LHC anomalies, as given in Eqs.
+(1)–(4), a lower value of $\rho_{t/\tau}^{S}$ compared to the parameter space shown in Fig 1 will be enough, which corroborates Iguro:2022dok .
+Note that, in Fig.
+1, the parameter space does not depend on the sign of the angle $\alpha$, and thus, the results are identical for both $\pm\alpha$.
 3 Uncovering the CP Nature of the Resonance
-The CP transformation property of a scalar particle can be inferred from its decay to tau pairs since the correlation of the tau lepton polarisation planes is governed by the CP properties of the parent particle.
-From the reconstruction of the decay planes of the tau leptons, it is possible to infer the CP properties of the scalar particle.
-The differential decay width of the scalar boson can be written as Kramer:1993jn ; Berge:2014sra ; Berge:2014wta
-$d\Gamma_{S\to\tau^{+}\tau^{-}}\propto 1-\dfrac{\pi^{2}}{16}b(E_{+})b(E_{-})\cos(\phi_{CP}-2\alpha),$ (10)
+The CP transformation property of a scalar particle can be inferred from its decay to tau pairs since the correlation of the tau lepton polarization planes is governed by the CP properties of the parent particle.
+From the reconstruction of the decay planes of the tau leptons, it is then possible to infer the CP properties of the scalar particle.
+3.1 Hadronic $\tau^{\pm}$ decays
+The differential decay width of a spin-0 boson into two fermions can be written as Kramer:1993jn ; Berge:2014sra ; Berge:2014wta
+$d\Gamma_{S\to\tau^{+}\tau^{-}}\propto 1-\dfrac{\pi^{2}}{16}b(E_{+})b(E_{-})\cos(\phi_{\rm CP}-2\alpha),$ (10)
 where $b(E_{\pm})$ are the spectral function describing the spin analyzing power of a given decay mode Tsai:1971vv ; Berge:2014wta .
-For $\tau\to\pi\nu$ decay, $b(E_{\pm})=1$, and for other decay channels, it depends on the energy of the charged hadron at their respective tau-lepton rest frame.
-There are several methods  Bower:2002zx ; Desch:2003rw ; Berge:2013jra ; Berge:2008wi ; Berge:2008dr ; Berge:2011ij ; Berge:2014sra ; Berge:2014wta ; Berge:2015nua to estimate $\phi_{CP}$ and in this work we have used Impact Parameter (IP) method Berge:2014sra ; Berge:2015nua and the $\rho$ decay plane method  Bower:2002zx ; Berge:2015nua ; Han:2016bvf ; Dutta:2021del for hadronic $\tau$ decay.
-For completeness, we have described the methods briefly.
-3.1 Impact Parameter Method
-In the Impact Parameter (IP) method Berge:2014sra ; Berge:2015nua , the $\tau$-lepton decay plane is constructed using the 3-momentum vector ($\vec{q}^{\pm}$) and the three-dimensional (3D) impact parameter vector ($\vec{n}_{\pm}$) of the charged particles.
-The method can be used for any one-prong decay of the tau lepton, and in this work we have implemented the method for $S\to\tau^{+}\tau^{-}\to(\pi^{+}\nu_{\tau})(\pi^{-}\bar{\nu}_{\tau})$ decay.
-The IP vector in the laboratory frame is defined as the distance of the closest approach of the charged particle’s ($\pi^{\pm}$) track to the reconstructed primary vertex (PV) of an event.
-It is possible to reconstruct the 3D IP vector using the distance between the point of closest approach and the primary vertex in the $x-y$-plane ($d_{0}$) and the $z$-coordinate of the point ($z_{0}$ansen:2020ecn .
-We have used $d_{0}$ and $z_{0}$ and charged track momentum (highest $p_{T}$ charged track inside a $\tau$-tagged jet) provided by Delphes to find out the IP vector using the procedure describe in Hansen:2020ecn .
-The IP vector then converted to a four vector $n_{\pm}^{\mu}=(0,\vec{n}_{\pm})^{T}$ and boosted to zero-momentum frame (ZMF), the $\pi^{+}\pi^{-}$ rest frame.
-The boosted IP four-vectors are denoted as $n_{\pm}^{*\mu}$ and we define the angle,
-$\phi^{*}=\arccos(\vec{n}^{*}_{+_{\perp}}.~{}\vec{n}^{*}_{-_{\perp}}),$ (11)
-where $n^{*\mu}_{\pm_{\perp}}$ is the perpendicular component of $n^{*\mu}_{\pm}$ along the charged track $h^{\pm}$.
-To obtain a signed angle, we define the triple product, $\mathcal{O}(=q^{*-}\cdot(\vec{n}^{*}_{+_{\perp}}.~{}\vec{n}^{*}_{-_{\perp}}))$ where $q^{*-}$ is the 3-momentum vector of $\pi^{-}$ in ZMF.
-The CP observable is given by,
-$\phi_{CP}=\left\{\begin{aligned} &\tilde{\phi}^{*}+\pi\quad&\mathcal{O}\geq 0\\ &\tilde{\phi}^{*}\quad&\mathcal{O}<0.\end{aligned}\right.$ (12)
-3.2 Rho Decay Plane Method
-When the $\tau$ lepton decays via the rho meson, $\tau\to\rho^{\pm}\nu_{\tau}\to\pi^{\pm}\pi^{0}\nu_{\tau}$, we use the information from the charged and neutral pion to define a CP-odd observable.
+For $\tau^{\pm}\to\pi^{\pm}\nu$ decays, $b(E_{\pm})=1$ while, for other decay channels, it depends on the energy of the charged hadron at the respective tau-lepton rest frame.
+There are several methods  Bower:2002zx ; Desch:2003rw ; Berge:2013jra ; Berge:2008wi ; Berge:2008dr ; Berge:2011ij ; Berge:2014sra ; Berge:2014wta ; Berge:2015nua to estimate $\phi_{\rm CP}$ and in this work we have used the Impact Parameter (IP) one Berge:2014sra ; Berge:2015nua and the $\rho$ decay plane one Bower:2002zx ; Berge:2015nua ; Han:2016bvf ; Dutta:2021del , for hadronic $\tau$ decays.
+For completeness, we will briefly describe the two methods next.
+3.1.1 Impact Parameter Method
+[…]
+$\phi^{*}=\arccos(\hat{n}^{*}_{+_{\perp}}.~\hat{n}^{*}_{-_{\perp}}),$ (11)
+where $\hat{n}^{*}_{\pm_{\perp}}$ is the normalized perpendicular component of $\vec{n}^{*}_{\pm}$ along the charged track $\pi^{\pm}$.
+To obtain a signed angle, we define the triple product, $\mathcal{O}(=\hat{q}^{*}_{-}\cdot(\hat{n}^{*}_{+_{\perp}}.~\hat{n}^{*}_{-_{\perp}}))$ where $\hat{q}^{*}_{-}$ is the normalized 3-momentum vector of $\pi^{-}$ in the ZMF.
+The CP observable is given by
+$\phi_{\rm CP}=\left\{\begin{aligned} &\phi^{*}\quad&\mathcal{O}\geq 0\\ &2\pi-\phi^{*}\quad&\mathcal{O}<0.\end{aligned}\right$. (12)
+3.1.2 Rho Decay Plane Method
+When the $\tau$ lepton decays via the rho meson, $\tau^{\pm}\to\rho^{\pm}\nu_{\tau}\to\pi^{\pm}\pi^{0}\nu_{\tau}$, we use the information from the charged and neutral pion to define a CP-odd observable.
 In the $\pi^{+}\pi^{-}$ rest frame, the acoplanarity angle $\phi^{*}$ is defined between the plane spanned by the $(\pi^{+},\pi^{0})$ and $(\pi^{-},\pi^{0})$ vectors.
 The acoplanarity angle is given by Bower:2002zx ; Berge:2015nua ; Hansen:2020ecn ,
-$\phi^{*}=\arccos(\vec{q}^{~{}*0}_{+_{\perp}}.~{}\vec{q}^{~{}*0}_{-_{\perp}}),$ (13)
-where $\vec{q}^{~{}*0}_{+(-)_{\perp}}$ is the 3-momentum of the $\pi^{0}$ associated with $\pi^{+(-)}$ in the $\pi^{+}\pi^{-}$ rest-frame transverse to the direction of the associated charged pion.
-The signed angle $\tilde{\phi}^{*}$ is obtained using the CP-odd triple correlation $\mathcal{O_{CP}}(=q^{*-}\cdot(q^{*0}_{+_{\perp}}\times~{}q^{*0}_{-_{\perp}})$) and we define
-$\tilde{\phi}^{*}=\left\{\begin{aligned} &\phi^{*}\quad&\mathcal{O_{CP}}\geq 0\\ &2\pi-\phi^{*}\quad&\mathcal{O_{CP}}<0.\end{aligned}\right.$ (14)
+$\phi^{*}=\arccos(\hat{q}^{~*0}_{+_{\perp}}.~\hat{q}^{~*0}_{-_{\perp}}),$ (13)
+where $\hat{q}^{~*0}_{\pm_{\perp}}$ is the normalized 3-momentum of the $\pi^{0}$ associated with $\pi^{\pm}$ in the $\pi^{+}\pi^{-}$ rest-frame transverse to the direction of the associated charged pion.
+The signed angle $\tilde{\phi}^{*}$ is obtained using the CP-odd triple correlation $\mathcal{O_{\rm CP}}(=\hat{q}^{*}_{-}\cdot(\hat{q}^{*0}_{+_{\perp}}\times~\hat{q}^{*0}_{-_{\perp}})$) and we define
+$\tilde{\phi}^{*}=\left\{\begin{aligned} &\phi^{*}\quad&\mathcal{O_{\rm CP}}\geq 0\\ &2\pi-\phi^{*}\quad&\mathcal{O_{\rm CP}}<0.\end{aligned}\right$. (14)
 In the $S\to\tau^{+}\tau^{-}\to(\pi^{+}\pi^{0}\nu_{\tau})(\pi^{-}\pi^{0}\bar{\nu}_{\tau})$ decay, the spin analyzing power is proportional to $y_{\rho}=\dfrac{(E_{\pi^{+}}-E_{\pi^{0}})}{(E_{\pi^{+}}+E_{\pi^{0}})}\times\dfrac{(E_{\pi^{-}}-E_{\pi^{0}})}{(E_{\pi^{-}}+E_{\pi^{0}})}$ at the respective $\tau^{\pm}$ rest frame Bower:2002zx .
-Since $y_{\rho}$ can be both positive and negative and if we integrate over the charged and neutral pion moments, $\phi_{CP}$ does not display $CP$ sensitivity.
-We categorize the events depending on $y_{\rho}$ and define the final $CP$ sensitive variable,
-$\phi_{CP}=\left\{\begin{aligned} &\tilde{\phi}^{*}+\pi\quad&y_{\rho}<0\\ &\tilde{\phi}^{*}\quad&\textrm{ otherwise}.\end{aligned}\right.$ (15)
+Since $y_{\rho}$ can be both positive and negative and if we integrate over the charged and neutral pion moments, $\phi_{\rm CP}$ does not display CP sensitivity.
+We categorize the events depending on $y_{\rho}$ and define the final CP-sensitive variable,
+$\phi_{\rm CP}=\left\{\begin{aligned} &\tilde{\phi}^{*}+\pi\quad&y_{\rho}<0\\ &\tilde{\phi}^{*}\quad&\textrm{ otherwise}.\end{aligned}\right$. (15)
 To estimate $y_{\rho}$, we have calculated the energies of the pions in the lab frame, as it is not possible to reconstruct the $\tau$ rest frames.
-The choice of lab frame energy does not significantly reduce the asymmetry in $\phi_{CP}$ as long as the $\rho$ mesons are energetic ($p_{T}(\rho^{\pm})>20$ GeV) Berge:2015nua which is in general true for LHC.
-3.3 Combined IP and Rho Decay Plane Methods
+The choice of lab frame energy does not significantly reduce the asymmetry in $\phi_{\rm CP}$ as long as the $\rho$ mesons are energetic ($p_{T}(\rho^{\pm})>20$ GeV) Berge:2015nua which is in general true for LHC.
+3.1.3 Combined IP and Rho Decay Plane Methods
 In the case of mixed decay, $\tau^{+}\tau^{-}\to(\pi\nu)(\rho\nu)$, we can combine the above-mentioned methods to define a CP-sensitive observable.
 This is possible if we consider the $\pi^{+}\pi^{-}$ rest frame for the $\rho$ Decay plane method.
-In the case of $\tau^{+}\tau^{-}\to(\rho^{+}\nu_{\tau})(\pi^{-}\bar{\nu}_{\tau})$ decay, the angle is defined by Berge:2015nua ,
-$\phi^{*}=\arccos(\vec{q}^{~{}*0}_{+_{\perp}}.~{}\vec{n}^{~{}*}_{-_{\perp}}),$ (16)
-where, $\vec{q}^{~{}*}_{\pm}$ is the 3-momentum of $\pi^{\pm}$ and $\vec{n}^{~{}*}_{-}$ is the IP vector of $\pi^{-}$ and tall the vectors are in the $\pi^{+}\pi^{-}$ rest frame.
-As before, we can define a signed angle depending on the value of the triple product $\mathcal{O}_{3}(=\vec{q}^{~{}*}_{-}\cdot(\vec{q}^{~{}*0}_{+_{\perp}}\times~{}\vec{n}^{~{}*}_{-_{\perp}})$).
-On the other hand, if $\tau^{+}$ decay to $\pi^{+}\nu_{\tau}$, we can define $\phi^{*}$ and the triple product appropriately.
-Note that, in both cases of mixed decay, we have considered the event classification based on $y_{\pm}(=(E_{\pi^{\pm}}-E_{\pi^{0}}))$.
+In the case of $\tau^{+}\tau^{-}\to(\rho^{+}\nu_{\tau})(\pi^{-}\bar{\nu}_{\tau})$ decays, the angle is defined by Berge:2015nua ,
+$\phi^{*}=\arccos(\hat{q}^{~*0}_{+_{\perp}}.~\hat{n}^{~*}_{-_{\perp}}),$ (16)
+where, $\hat{q}^{~*0}_{+}$ is the normalized 3-momentum of $\pi^{0}$ (originating from $\rho^{+}$) and $\hat{n}^{~*}_{-}$ is the IP vector of $\pi^{-}$ and all the vectors are in the $\pi^{+}\pi^{-}$ rest frame.
+As before, we can define a signed angle depending on the value of the triple product $\mathcal{O}_{3}(=\hat{q}^{~*}_{-}\cdot(\hat{q}^{~*0}_{+_{\perp}}\times~\hat{n}^{~*}_{-_{\perp}})$).
+Instead, if $\tau^{+}$ decays to $\pi^{+}\nu_{\tau}$, we can define $\phi^{*}$ and the triple product appropriately.
+Note that, in both cases of mixed decay, we have considered the event classification based on $y^{\rho}_{\pm}\Big{(}=\dfrac{E_{\pi^{\pm}}-E_{\pi^{0}}}{E_{\pi^{\pm}}+E_{\pi^{0}}}\Big{)}$.
+3.2 Semi-leptonic Decays
+For the semi-leptonic channels, $S\to\tau^{+}\tau^{-}\to(\ell^{+}\nu_{\ell}\bar{\nu}\tau)(\tau_{h}^{-}\nu\tau)$ (and charge conjugate), where the leptons denote electrons and muons, the hadronic $\tau$ decay is reconstructed using either the IP method or the $\rho$ decay plane method, as described earlier.
+The leptonic side is treated as a one-prong decay, and we are using the impact parameter of the charged lepton for our analysis.
+[…]
 3.3 Event Simulation and Selection
@@
 To handle the $\tau$ polarization, we have used the TauDecay Hagiwara:2012vz package in MadGraph5-aMC@NLO-3.5.3.
-For showering and hadronization we used PYTHIA-8.3 Sjostrand:2014zea ; Bierlich:2022pfr , and detector effects were simulated using Delphes-3.5.0deFavereau:2013fsa .
-We use anti-kt algorithm Cacciari:2008gp for jet clustering with radius parameter $R=0.5$ and $p_{T}(j)>20$ GeV.
-For $\tau$-tagging we have used medium tag point with $70\%$ tagging and $5\times 10^{-3}$ mis-tagging efficiency.
-The signal events are classified as two $\tau$-tagged jets with $p_{T}>40$ GeV and $|\eta|<2.1$.
-We veto events with any $b$-tagged jets.
-The background for such signal at the LHC is dominated by $Z\to\tau^{+}\tau^{-}$, $W^{\pm}+jets$ and QCD multi-jet events.
-We have generated sample background data to understand the nature of BG distribution for $CP$ sensitive observable.
-As we will discuss later, we have used actual data provided by the CMS collaboration through HEPDATA Maguire:2017ypu and rescaled it accordingly for different luminosity.
+For showering and hadronization we used PYTHIA-8.3 Sjostrand:2014zea ; Bierlich:2022pfr while detector effects were simulated using Delphes-3.5.0 deFavereau:2013fsa .
+We use the anti-kt algorithm Cacciari:2008gp for jet clustering with radius parameter $R=0.5$ and $p_{T}(j)>20$ GeV.
+For $\tau$-tagging, we have used a medium tag point with $70\%$ tagging and $5\times 10^{-3}$ mis-tagging efficiency.
+We follow Ref.
+CMS:2022goy for event selection.
+For the hadronic channel, the signal events are classified as two $\tau$-tagged jets with $p_{T}>40$ GeV and $|\eta|<2.1$.
+For the semi-leptonic events, electrons (muons) should have $|\eta|<2.1$, and need to satisfy isolation criterion of $I_{\text{rel}}<0.15$.
+The electrons (muons) with $p_{T}>$ 25 GeV (20) GeV is required for the $e-\tau_{h}$ ($\mu-\tau_{h}$) mode.
+The hadronically decaying tau ($\tau_{h}$) candidate is required to have $|\eta|<2.3$.
+A $p_{T}$ threshold of 35 GeV (32 GeV) is applied when the event is selected by an $e-\tau_{h}$ ($\mu-\tau_{h}$) pair modes.
+For both the hadronic and the semi-leptonoc modes, We veto events with any $b$-tagged jets, and we have considered events having $p_{T}^{\tau^{+}\tau^{-}}>200$ GeV as shown in the right panel of Fig.
+9 in Ref.
+CMS:2022goy , where $p_{T}^{\tau^{+}\tau^{-}}$ is obtained by the vectorial $p_{T}$ sum of the visible decay products of tau leptons.
+Such $p_{T}^{\tau^{+}\tau^{-}}$ criteria help to substantially minimize the background events.
 We obtain the charged and neutral pions from the $\tau$-jet constituents.
 From all the tracks within each $\tau$-jet, the track with the highest $p_{T}$ is considered to be the charged pion.
-Neutral pions decay promptly into a pair of photons and therefore deposit their energies in the electromagnetic calorimeter, ECAL.
+Neutral pions decay promptly into a pair of photons and therefore deposit their energies in the electromagnetic calorimeter.
 From all the energy deposits, called towers, we estimate the one with the highest $E_{T}$ and assign the tower to the neutral pion.
-For the combination method, the charged pion track with higher $p_{T}$ is associated with the two-body decay of the $\tau$ ($\tau\to\pi\nu_{\tau}$).
+For the IP-$\rho$ combination method, the charged pion track with higher $p_{T}$ is associated with the two-body decay of the $\tau^{\pm}$ ($\tau^{\pm}\to\pi^{\pm}\nu_{\tau}$).
+In the case of the semi-leptonic mode, the track associated with the lepton is considered while, for the hadronic part, we have used either the IP or the $\rho$-decay method.
+The background for such signal at the LHC is dominated by $Z\to\tau^{+}\tau^{-}$, $W$ + jets, $t\bar{t}$ and QCD multi-jet events.
+We have generated sample background data to understand the nature of the corresponding distribution for our CP-sensitive observable.
+As we will discuss later, we have used actual data provided by the CMS collaboration through HEPDATA Maguire:2017ypu ; hepdata.128147 and rescaled the background accordingly for a different luminosity.
 4 Results
-[…]
-Now, we show how to uncover the CP properties of the scalar particle at the LHC.
+Figure 2: Signal and background distribution of our CP-sensitive observable for combining all the hadronic (left panels) and semi-leptonic (right panels) modes with an integrated luminosity of 3000 fb-1.
+The top panel shows the distribution for a CP-even 95 GeV scalar, while the middle panel shows the same for a maximally CP-mixed state, and the CP-odd distribution is shown in the bottom panel.
+Figure 3: Signal and theoretical distribution fo the CP observable $\phi_{\rm CP}$ scaled to HL-LHC from the observed data at 139 fb-1.
+The top panels show the distribution for CP-even (top-left) and maximally CP-mixed (top-right) scalar, and the bottom panels show the distribution for a CP-odd state.
+The hadronic IP method provides the best reconstruction and closely resembles the theoretical distribution, whereas the other methods show the exact nature of the distribution with subdued extrema.
+If $\alpha$ becomes $\pi/4$, then the peak of the distribution will be at $\phi_{\rm CP}=\pi/2$.
+Now, we show how to access the CP properties of the spin-0 with particle with mass around 95 GeV at the LHC.
 For a realistic representation of signal and background events, we have used the present CMS data of the $\tau^{+}\tau^{-}$ excess CMS:2018cyk ; CMS:2023yay .
-The data is taken from HEPDATAMaguire:2017ypu ; hepdata.128147 .
-For our analysis, we have selected the events for which the invariant mass $m_{\tau^{+}\tau^{-}}$ lies within $60-120$ GeV and we have selected the events with $p_{T}^{\tau^{+}\tau^{-}}>200$ GeV obtained by the vectorial $p_{T}$ sum of the visible decay products of tau lepton.
-This helps to minimise the background events significantly, as evident from Fig.
-9 in ref CMS:2022goy .
-To demonstrate the CP nature of the scalar, we plot the $\phi_{CP}$ for both signal and background events.
-The distribution peaks at a particular value of $\phi_{CP}$ for a given CP-admixture, whereas the background remains primarily flat.
-This enables estimating the CP nature from the $\phi_{CP}$ distribution alone.
-We found that with an integrated luminosity of 138 $fb^{-1}$, the total event distribution remains within statistical uncertainty of the background ($\sqrt{N_{B}}$).
-We considered high luminosity LHC (HL-LHC) Gianotti:2002xx ; Apollinari:2015wtw , and appropriately rescaled the existing data to an integrated luminosity of 3000 $fb^{-1}$.
-The $\phi_{CP}$ distribution for signal and background events for HL-LHC is shown in Fig.
-2, where background events are shown in yellow with statistical uncertainty bars, and the signal events are shown in green.
-As anticipated, the background events exhibit a near-uniform distribution in all the cases.
-The top panel depicts the distribution for CP-Even (top-left) and CP-odd (top-right) scalar, whereas the bottom panels depict maximally CP-Mixed states.
-From all the figures, it is evident that for a given CP nature of the resonance, the cumulative number of signal and background events rise beyond the background uncertainty, signalling the CP property of the scalar.
-$\alpha$ $\phi_{CP}$ Window $\mathcal{S}_{0}$ $\mathcal{S}_{\sqrt{b}}$ $\mathcal{S}_{1\%}$ $\mathcal{S}_{10\%}$
-00 $120^{\circ}-240^{\circ}$ $3.10$ $2.18$ $2.01$ $0.26$
-$\pi/2$ $0^{\circ}-80^{\circ}~{}~{}\&~{}~{}280^{\circ}-360^{\circ}$ $3.51$ $2.47$ $2.09$ $0.26$
-$\pi/4$ $200^{\circ}-320^{\circ}$ $3.13$ $2.2$ $2.03$ $0.27$
--$\pi/4$ $40^{\circ}-160^{\circ}$ $3.12$ $2.2$ $2.03$ $0.27$
-Table 1: Significance of different CP eigenstates with various background uncertainties at HL-LHC.
-In order to quantify the result, we have chosen a window in $\phi_{CP}$ distribution and estimated a signal significance using the following formula Cowan:2010js ,
-$\mathcal{S}_{\sigma_{b}}=\sqrt{2}\Big{\{}(N_{S}+N_{B})\ln\Big{[}\frac{(N_{S}+N_{B})(N_{B}+\sigma_{b}^{2})}{N_{B}^{2}+(N_{S}+N_{B})\sigma_{b}^{2}}\Big{]}-\frac{N_{B}^{2}}{\sigma_{b}^{2}}\ln\Big{[}1+\frac{\sigma^{2}_{B}N_{S}}{N_{B}(N_{B}+\sigma_{b}^{2})}\Big{]}\Big{\}}^{1/2},$ (17)
-where $N_{S(B)}$ denotes signal(background) events and $\sigma_{b}$ denotes the background uncertainty.
-In Tab.
-1 we have tabulated the significance for an optimal $\phi_{CP}$ window with various background uncertainty to achieve a robust estimation of the Significance.
-As expected, vanishing BG uncertainty provides a significance of more than $3~{}\sigma$ at HL-LHC.
-As uncertainty increases, Significance reduces but remains large enough if we consider the background uncertainty solely statistical ($\sqrt{N_{B}}$) or $1\%$.
-If $\sigma_{b}$ becomes large ($\sim 10\%$), the number of signal events need to be higher to achieve any significance.
+The data is taken from HEPDATA and for our analysis, we have selected the events for which the invariant mass $m_{\tau^{+}\tau^{-}}$ lies within 6012060-120 GeV with $p_{T}^{\tau^{+}\tau^{-}}>200$ GeV.
+To demonstrate the CP nature of the spin-0 object, we obtain the $\phi_{\rm CP}$ distribution for both signal and background events.
+We found that, with an integrated luminosity of 138 fb-1, the total event distribution remains within statistical uncertainty of the background ($\sqrt{N_{B}}$) and that the situation is not considerably better at the end of Run 3.
+Hence, we considered the HL-LHC Gianotti:2002xx ; Apollinari:2015wtw and appropriately rescaled the existing data to an integrated luminosity of 3000 fb-1.
+The $\phi_{\rm CP}$ distribution for signal and background events for the HL-LHC is shown in Fig.
+2, where background events are shown in yellow, whereas the signal events are shown in green.
+The combined hadronic events are shown in the left panel, whereas the right panel depicts semi-leptonic events after combining all the modes.
+The top, middle and bottom panels in Fig.
+2 depicts the distribution for a CP-even, maximally CP-mixed and CP-odd state, respectively.
+As anticipated, the background events exhibit a near-uniform distribution in all cases.
+As theoretically expected, the signal distribution peaks at $\phi_{\rm CP}=\pi(0\textrm{ and }2\pi)$ for a CP-even(odd) spin-0 state, whereas for the CP-mixed one, the peak is at $3\pi/2$.
+From all the figures, it is evident that, for a given CP nature of the resonance, the cumulative number of signal events rises beyond the background uncertainty, thus signalling the CP properties of the new particle.
+Figure 4: Plot of $\Delta\chi^{2}$ as a function of fitting angle $\alpha_{fit}$ assuming that the 95 GeV scalar to be CP-even(red), maximal CP-mixed(blue) and CP-odd(green).
+The solid(dashed) line corresponds to background systematics assumed to be 0.5%(1%).
+The black horizontal line at $\Delta\chi^{2}=13.36$ shows the 90% Confidence Level (CL) interval.
+To quantify our results, we use the $\chi^{2}$ defined as
+$\chi^{2}(\alpha)=\sum_{\rm Modes}\quad\sum_{i\in\rm{Bins}}\dfrac{\left(S_{i}^{\alpha_{H}}-\frac{n_{S}}{\Gamma}\frac{d\Gamma}{d\phi_{\rm CP}}(\alpha)\right)^{2}}{(\delta S_{i})^{2}+(\delta_{sys}^{2})},$ (17)
+where $\alpha_{H}$ is the assumed hypothesis (i.e., the CP nature) for the 95 GeV state and $S_{i}^{\alpha_{H}}$ denotes the signal events in the $i^{th}$ bin.
+The input parameter $\alpha$ provides the theoretical estimation of the CP observable $\phi_{CP}$ normalized to the total number of signal events, $n_{S}$.
+For each bin, the variance includes the statistical component defined as $\delta S_{i}=\sqrt{S_{i}}$ and an uncorrelated systematic uncertainty corresponding to 0.5% or 1% of the expected background in that bin.
+Finally, the summation over modes includes all the semi-leptonic and hadronic $\tau^{+}\tau^{-}$ divided into IP, $\rho$ and combined IP-$\rho$ modes.
+The normalized binned distribution of signal events $(S_{i}^{\alpha_{H}})$ as well as the theoretical distribution $\left(\dfrac{n_{S}}{\Gamma}\dfrac{d\Gamma}{d\phi_{\rm CP}}(\alpha)\right)$ for various channels is shown in Fig.
+3 for a CP-even (top-left), maximally mixed-CP state (top-right) as well as CP-odd (bottom).
+As expected, the hadronic IP method yields the most accurate distribution, closely matching the theoretical distribution, while the other methods capture the overall shape, but with maxima and minima systematically suppressed relative to the theoretical distribution.
+In Fig.
+4, we present our main result: the distribution of $\Delta\chi^{2}$ as a function of $\alpha_{\text{fit}}$.
+The red, blue and green lines correspond to the assumed hypotheses that the 95 GeV scalar is CP-even, maximally CP-mixed, and CP-odd, respectively.
+The solid (dashed) curves represent $\delta_{\text{sys}}=0.5\%\ (1\%)$ of the total background, which enters the $\chi^{2}$ calculation as defined in Eq.
+(17).
+The signal distribution is binned with $N_{\text{bin}}=9$ and the $90\%$ CL interval corresponds to $\Delta\chi^{2}=13.36$, shown as a black horizontal line in Fig.
+4.
+We find that, at the HL-LHC, it will be possible to ascertain the CP nature of the 95 GeV scalar within $\pm 0.27\ (0.47)$ radians of the true hypothesis at $90\%$ CL, depending on the assumed background systematics.
 5 Conclusions
-In summary, under the assumption that several anomalies seen in the invariant mass of $\gamma\gamma$, $\tau^{+}\tau^{-}$ (from LHC data) and $b\bar{b}$ (from LEP data) events are due to a single spin-0 object with mass 95 GeV having Yukawa couplings to third generation fermions, we have performed a detector level analysis aimed at extracting the CP properties of such an object.
-We have done so limitedly to the $\tau^{+}\tau^{-}$ final state, as it offers the possibility of accessing the corresponding CP quantum numbers whichever the $\tau$ decay, whether leptonic or hadronic, through the relative angle between the two decay planes defined through the visible particles emerging from the $\tau$ lepton decays.
+In summary, under the assumption that several anomalies seen in the invariant mass of $\gamma\gamma$, $\tau^{+}\tau^{-}$ (from LHC data) and $b\bar{b}$ (from LEP data) events are due to a single spin-0 object with mass approximately 95 GeV having Yukawa couplings to third generation fermions, we have performed a detector level analysis aimed at extracting the CP properties of such an object.
+We have done so limitedly to the $\tau^{+}\tau^{-}$ final state, as it offers the possibility of accessing the corresponding CP quantum numbers, whichever the $\tau$ decay, whether leptonic or hadronic, through the relative angle between the two decay planes defined through the visible particles emerging from the $\tau$ lepton decays.
 Our approach has been based on a simplified model, which is characterized by the relevant Yukawa interactions being defined in terms of correcting factors to the corresponding SM vertex strengths while allowing for both a scalar and pseudoscalar interaction as well as any superposition thereof.
 We have found that, under the assumption that the $\gamma\gamma$ and $\tau^{+}\tau^{-}$ anomalies will persist throughout the LHC era, the HL-LHC, in standard energy and luminosity configurations, will offer the chance to extract the CP properties of such a new spin-0 resonance.
-Once the latter is combined with cross-section and BR information, it would then be possible to map the results obtained in our model-independent parameterization into well-defined models like, e.g., a 2HDM of whichever Type, so long that it can host a suitable 95 GeV state in its Higgs particle spectrum, as it has indeed been proven to be the case in previous literature.
+Once the latter is combined with cross section and BR information, it would then be possible to map the results obtained in our model-independent parameterization into well-defined models like, e.g., a 2HDM of whichever Type, so long that the latter can host a suitable 95 GeV state in its Higgs particle spectrum.
 We therefore advocate experimental analyses by ATLAS and CMS aimed at implementing our approach, or indeed improved versions of it, with upcoming datasets at the CERN machine.
@@
 ST/L000296/1.
+P.S.
+is supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education through the Center for Quantum Spacetime (CQUeST) of Sogang University (RS-2020-NR049598).
+P.S.
+would also like to thank the KIAS Center for Advanced Computation for providing computing resources.